OpenStudy (anonymous):

solve each equation. Check your answers and identify any extraneous roots. x - 2/x^2 - x - 6 = 1/x^2-4 + 3/2x + 4=

OpenStudy (anonymous):

@ganeshie8 think u can help?

ganeshie8 (ganeshie8):

is the question pasted correctly

ganeshie8 (ganeshie8):

= in the end is throwing me off

OpenStudy (anonymous):

ignore that equal sign.

OpenStudy (anonymous):

i had to write it out cause its on paper. -.-

ganeshie8 (ganeshie8):

\(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4}\)

ganeshie8 (ganeshie8):

in your paper, is the question looking as above ?

OpenStudy (anonymous):

yess

ganeshie8 (ganeshie8):

this requires lot of paper !

OpenStudy (anonymous):

ya it looks like it..lol

ganeshie8 (ganeshie8):

ok lets do it

ganeshie8 (ganeshie8):

do you knw factoring ?

OpenStudy (anonymous):

not really i never took algebra 2 in a classroom so i dont know anything :/

ganeshie8 (ganeshie8):

its okay factoring is easy

ganeshie8 (ganeshie8):

lets start wid factoring bottom :- x^2-x-6

ganeshie8 (ganeshie8):

we need to write that as product of two things. thats called factoring

OpenStudy (anonymous):

oh ok

ganeshie8 (ganeshie8):

this is how we going to factor it : x^2 - x - 6 x^2 -3x + 2x -6 x(x-3) + 2(x-3) (x-3)(x+2)

ganeshie8 (ganeshie8):

make sure you understand how we factored, we going to do the same for many future problems. so you need to understand them if u want to feel comfortable when solving other problems later. ok. so just make sure u understand how to factor. if u dont, just ask ok

ganeshie8 (ganeshie8):

take ur time

OpenStudy (anonymous):

ok how did you get 3x & the extra 2x?

ganeshie8 (ganeshie8):

good q, didnt u notice x is missing

ganeshie8 (ganeshie8):

x^2 - x - 6 x^2 -3x + 2x -6

OpenStudy (anonymous):

yes & also the addition symbol

ganeshie8 (ganeshie8):

we need to rewrite the middle "-x"

ganeshie8 (ganeshie8):

i wrote, "-3x + 2x" in place of "-x"

ganeshie8 (ganeshie8):

il tell why we picked those soon, but do u agree that there is nothing wrong in what we did ?

ganeshie8 (ganeshie8):

cuz, -3x + 2x = -x

OpenStudy (anonymous):

yes i seee

ganeshie8 (ganeshie8):

ok, lets move on. (il tell you how to pick the numbers later)

ganeshie8 (ganeshie8):

x^2 - x - 6 x^2 -3x + 2x -6

OpenStudy (anonymous):

ok sounds good!

ganeshie8 (ganeshie8):

from the first two terms, take x common

ganeshie8 (ganeshie8):

x^2 - x - 6 x^2 -3x + 2x -6 x(x-3) + 2x - 6

ganeshie8 (ganeshie8):

still wid me ?

ganeshie8 (ganeshie8):

from the last two terms take 2 common

ganeshie8 (ganeshie8):

x^2 - x - 6 x^2 -3x + 2x -6 x(x-3) + 2x - 6 x(x-3) + 2(x-3)

ganeshie8 (ganeshie8):

now we see (x-3) is common in both terms, so take it common !

ganeshie8 (ganeshie8):

x^2 - x - 6 x^2 -3x + 2x -6 x(x-3) + 2x - 6 x(x-3) + 2(x-3) (x-3)(x+2)

OpenStudy (anonymous):

ok im kinda with you :p

ganeshie8 (ganeshie8):

we are done factoring ! just go thru and make yourself convinced u understand

ganeshie8 (ganeshie8):

so bottom x^2-x-6 can be written in factored form as (x-3)(x+2). lets go to our main equation

OpenStudy (anonymous):

okaay

ganeshie8 (ganeshie8):

\(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \)

ganeshie8 (ganeshie8):

\(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \)

ganeshie8 (ganeshie8):

next look at right side of equation. lets keep focusing on bottom

ganeshie8 (ganeshie8):

we have x^2-4

ganeshie8 (ganeshie8):

i want to factor it.

ganeshie8 (ganeshie8):

can u help me if u have any ideas how to factor it ?

OpenStudy (anonymous):

lemme try

ganeshie8 (ganeshie8):

heres the clue :- write 4 as 2^2 and use the formula, a^2-b^2 = (a+b)(a-b)

ganeshie8 (ganeshie8):

ok il help you wid this

ganeshie8 (ganeshie8):

x^2 - 4 x^2 - 2^2 (x+2)(x-2)

ganeshie8 (ganeshie8):

we used a^2-b^2 formula, when moving from 2nd step to 3rd step

ganeshie8 (ganeshie8):

we have our main working as :- \(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2x + 4} \)

ganeshie8 (ganeshie8):

look at the bottom of last term. 2x+4 2(x+2)

ganeshie8 (ganeshie8):

\(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2(x + 2)} \)

ganeshie8 (ganeshie8):

if you see (x+2) is there in the bottom for each term. so it can be cancelled

ganeshie8 (ganeshie8):

\(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2(x + 2)} \) \(\large \frac{ x - 2}{(x-3)\cancel{(x+2)}} = \frac{1}{\cancel{(x+2)}(x-2)} + \frac{3}{2\cancel{(x + 2)}} \) \(\large \frac{ x - 2}{(x-3)} = \frac{1}{(x-2)} + \frac{3}{2} \)

ganeshie8 (ganeshie8):

you still wid me or i lost u lol, this is a very lengthy prob. but simple one.

OpenStudy (anonymous):

ya im with you im just trying to write it out too to understand it >.<

ganeshie8 (ganeshie8):

good :) thats the only simplification possible by factoring. now we need to do the donkey work. get ready

OpenStudy (anonymous):

lol ok lets do this.

ganeshie8 (ganeshie8):

\(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2(x + 2)} \) \(\large \frac{ x - 2}{(x-3)\cancel{(x+2)}} = \frac{1}{\cancel{(x+2)}(x-2)} + \frac{3}{2\cancel{(x + 2)}} \) \(\large \frac{ x - 2}{(x-3)} = \frac{1}{(x-2)} + \frac{3}{2} \) \(\large \frac{ x - 2}{(x-3)} = \frac{2+3(x-2)}{2(x-2)}\) \(\large \frac{ x - 2}{(x-3)} = \frac{2+3x-6)}{2(x-2)}\) \(\large \frac{ x - 2}{(x-3)} = \frac{3x-4)}{2(x-2)}\)

ganeshie8 (ganeshie8):

i just took the LCD of right side, so we can write it as single term. instead of 2 terms. u knw LCD ?

OpenStudy (anonymous):

least common denominator?

ganeshie8 (ganeshie8):

exactly ! LCD of 2, (x-2) is simply 2(x-2)

ganeshie8 (ganeshie8):

let me knw once u follow upto above steps. we can go ahead with rest of the problem, then

OpenStudy (anonymous):

i follow

ganeshie8 (ganeshie8):

great :)

ganeshie8 (ganeshie8):

\(\large \frac{ x - 2}{x^2 - x - 6} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{x^2-4} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2x + 4} \) \(\large \frac{ x - 2}{(x-3)(x+2)} = \frac{1}{(x+2)(x-2)} + \frac{3}{2(x + 2)} \) \(\large \frac{ x - 2}{(x-3)\cancel{(x+2)}} = \frac{1}{\cancel{(x+2)}(x-2)} + \frac{3}{2\cancel{(x + 2)}} \) \(\large \frac{ x - 2}{(x-3)} = \frac{1}{(x-2)} + \frac{3}{2} \) \(\large \frac{ x - 2}{(x-3)} = \frac{2+3(x-2)}{2(x-2)}\) \(\large \frac{ x - 2}{(x-3)} = \frac{2+3x-6)}{2(x-2)}\) \(\large \frac{ x - 2}{(x-3)} = \frac{3x-4)}{2(x-2)}\) Cross multuply \(\large 2(x-2)^2 = (3x-4)(x-3)\)

OpenStudy (anonymous):

how do i set that up to cross multiply?

ganeshie8 (ganeshie8):